# Systems with Switching Causal Relations: A Meta-Causal Perspective

### v1

@@ 4 Applications
In this section, we discuss several applications of the meta-causal formalism. First, we revisit the motivational example and consider how meta-causal models can be used to attribute responsibility. Next, we identify the presence of multiple mechanisms for the bivariate case from sets of unlabeled data. Finally, we analyze the emergence of meta-causal states from a dynamical system, highlighting that meta-causal states can be more than simply conditioning the adjacency matrix of graphs on a latent variable. The code used in the experiments is available at https://anonymous.4open.science/r/metaCausal-CRL/.
@@ 4.1 Attributing Responsibility
Note that in this scenario, a classic counterfactual consideration, $A_{X}^{A_{\operatorname{\bm{\pi}}}=\text{standing still}}-A_{X}^{A_{\operatorname{\bm{\pi}}}=\text{following}}$, would also infer an effect of $A_{\operatorname{\bm{\pi}}}$ on $A_{X}$ using a purely value-based perspective. Attributing effects via observed changes in variable values is a valid approach, but fails to explain preventative mechanisms. For example, consider the simple scenario where two locks prevent a door from opening. Classical counterfactual analysis would attribute zero effect to the opening of either lock. A meta-causal perspective would reveal that the opening of the first lock already changes the meta-causal state of the model by removing its causal mechanisms that conditions the state of the door. In a sense, our meta-causal causal perspective is similar to that of the actual causality framework (Halpern, 2016; Chockler & Halpern, 2004). However, actual causality operates at the ‘actual’, i.e. value-based, level and does not take the mechanistic meta-causal view into perspective. Meta-causal models already allow us to reason about causal effects after opening the first lock due to a change in the causal graph, without us having to consider the effects of opening the second lock at some point to reach that conclusion.
@@ 4.2 Discovering Meta-Causal States in the Bivariate Case
Our goal in this experiment is to recover the number of meta-causal states $K\in[1..4]$ from data that exists between two variables $X,Y$ that are directly connected by a linear equation with added noise. We assume that each meta-causal state gives rise to a different linear equation $f_{k}:=\alpha_{k}X+\beta_{k}+\mathcal{N},k\in\mathbb{N}$, where $\alpha_{k},\beta_{k}$ are the slope and intercept of the respective mechanism and $\mathcal{N}$ is a zero-centered, symmetric, and quasiconvex noise distribution [^2] Implying unimodality and monotonic decreasing from zero allows us to distinguish the noise mean and intercept and to recover the parameters from a simple linear regression. Without loss of generality, we assume Laplacian noise, for which an L1-regression can estimate the true parameters of the linear equations (Hoyer et al., 2008). The causal direction of the mechanism is randomly chosen between different meta-causal states. The exact sampling parameters and plots of the resulting distributions are given in Sec. C and Fig. 3 in the appendix, respectively. In general, this scenario corresponds to the setting described above of inferring the values of a latent variable $Z$ (with $K=|Z|$) that indicate a particular meta-causal state of the system. Our goal is to recover the number and parameterizations of the causal mechanisms, and as a consequence to be able to reason about which points were generated by which mechanism.
@@
Approach. The problem we are trying to solve is twofold : first, we are initially unaware of the underlying meta-causal state $t$ that generated a particular data point $(x_{i},y_{i})$, which prevents us from estimating the parameterization $(\alpha_{k},\beta_{k})$ of the mechanism. Conversely, our lack of knowledge about the mechanism parameterizations $(\alpha_{k},\beta_{k})_{k\in[1..K]}$ prevents us from assigning class probabilities to the individual data points. Since neither the state assignment nor the mechanism parameterizations are initially known, we perform an Expectation-Maximization (EM; Dempster et al. (1977)) procedure to iteratively estimate and assign the observed data points to the discovered MCS parameterizations. Due to the local convergence properties of the EM algorithm, we further embed it into a locally optimized local random sample consensus approach (LO-RANSAC; Fischler & Bolles (1981); Chum et al. (2003)). RANSAC approaches repeatedly sample initial parameter configurations to avoid local minima, and successively perform several steps of local optimization - here the EM algorithm - to regress the true parameters of the mechanism.
@@
Assuming for the moment that the correct number of mechanisms $k^{*}$ has been chosen, we assume that the EM algorithm is able to regress the parameters of the mechanisms, $\alpha_{k}^{*},\beta_{k}^{*}$, whenever there exists a pair of points for each of the mechanisms, where both points of the pair are samples generated by that particular mechanism. The chances of sampling such an initial configuration decrease rapidly with an increasing number of mechanisms (e.g. $0.036\%$ probability for $k=4$ and equal class probabilities). Furthermore, we assume that the sampling probabilities of the individual mechanisms in the data can deviate from the mean by up to a certain factor $d$. In our experiments, we consider setups with $d\in\{0.0,0.1,0.2\}$. Given the number of classes and the maximum sample deviation of the mechanisms, one can compute an upper bound on the number of resamples required to have a $95\%$ chance of drawing at least one valid initialization. The bound is maximized by assigning the first half of the classes the maximum deviation probability $P_{\text{k-max}}=(1+d)/k$ and the other half the minimum deviation probability $P_{\text{k-min}}=(1-d)/k$. We provide the formulas for the upper bound estimation and in Sec. A in the Appendix and provide the calculated required resample counts in Table 2. In the worst case, for a scenario with $k=4$ mechanisms and $d=0.2$ maximum class probability deviation, nearly $14,900$ restarts of the EM algorithm are required, drastically increasing the potential runtime.
@@
1 Mechanism 2 Mechanisms 3 Mechanisms 4 Mechanisms
Max. Class Deviation = 0.0
Theoretical 1 23 363 8,179
Empirical 2 8 24 173
Max. Class Deviation = 0.1
Theoretical 1 26 429 10,659
Empirical 2 8 25 177
Max. Class Deviation = 0.2
Theoretical 1 30 526 14,859
Empirical 2 8 26 177
Table 2: Estimated Number of Required Resamples for Obtaining a 95% Convergence Rate with the LO-RANSAC Algorithm per Number of Mechanisms and Maximum Class Deviation. The empirical observed convergence rates of the EM algorithm drastically reduce the required number of samples of the theoretical derived bound. This reveals that our made assumptions where chosen to be quite conservative, and attests a good fit of the EM algorithm for regressing the mechanisms’ parameters.
@@ 1 Mechanism 2 Mechanisms 3 Mechanisms 4 Mechanisms
The EM algorithms return the estimated parameters $\alpha^{\prime}_{k},\beta^{\prime}_{k}$, and the mean standard deviation $b^{\prime}_{k}$ of the Laplacian [^3] Since mechanisms can go in both directions, $X\rightarrow Y$ and $Y\rightarrow X$, we repeat the regression for both directions and use an Anderson-Darling test (Anderson & Darling, 1952) on the residual to test which of the distributions more closely resembles a Laplacian distribution at each step.. This allows one to compute the class densities $k(x_{i},y_{i};\alpha^{\prime}_{k},\beta^{\prime}_{k},b^{\prime}_{k})$ for a pair of values $(x_{i},y_{i})$. Since the assignment of mechanisms to a data point may be ambiguous due to the overlap in the estimated PDFs, we normalize the density values of all mechanisms per data point and consider only those points for which the probability of the dominant mechanism is $0.4$ higher than the second class: $(\mathbf{x}^{j},\mathbf{y}^{j}):=((x_{i},y_{i})|\#1=j;f^{\#2}_{i}<P^{\#1}_{i}\times(1-0.4))$, where #$\#n$ indicates the class with the n-th highest density value. Finally, the residuals $\mathbf{y}^{j}-f^{\prime j}_{k}(\mathbf{x}^{j},\mathbf{y}^{j};\alpha^{\prime}_{k},\beta^{\prime}_{k},b^{\prime}_{k})$ are computed for all data points $(\mathbf{x}^{j},\mathbf{x}^{j})$ assigned to a particular predicted mechanism $f^{\prime j}_{k}$. We choose the parameterization that best fits the data for each $k\in[1..4]$ and, make use of the Anderson-Darling test (Anderson & Darling, 1952) to test the empirical distribution function of the residuals against the Laplacian distribution with an $\alpha=0.95$ using significance values estimated by Chen (2002). If all residual distributions of all mechanisms for a given $k^{\prime}$ pass the Anderson-Darling test, we choose that number as our predicted number of mechanisms. If the algorithm finds that none of the $k\in[1..4]$ setups pass the test, we refrain from making a decision. We provide the pseudo code for our method in Algorithm. 1 in the Appendix.
@@
Evaluation and Results. We evaluate our approach over all $k\in[1..4]$ by generating 100 different datasets for every particular number of mechanisms. For every data set we sample 500 data points from each mechanism $(x^{k}_{i},y^{k}_{i})=f_{k}(\alpha_{k}x^{k}_{i}+\beta_{k}+l_{i})$ where $l_{i}\sim\mathcal{L}(0,b_{k})$, using the same sampling method as before (c.f. Appendix C). Finally, we let the algorithm recover the number of mechanisms that generated the data.
@@
We compare theoretically computed and empirical in convergence results in Table 2. In practice, we observe that the convergence of the EM algorithm is more favorable than estimated, reducing from $23$ to $8$ required examples for the simple case of $k=2,d=0.0$, and requiring up to 83-times fewer samples for the most challenging setup of $k=4,d=0.2$, reducing from a theoretical of 14,859 to an empirical estimate of 177 samples. The actual convergence probabilities and the formula for deriving to sample counts are given in Table 3 and Sec. B of the Appendix.
@@
Extension to Meta-Causal State Discovery on Graphs. Our results indicate that identifying meta-causal mechanisms even in the bivariate case comes with an increasing number of uncertainty when it comes to increasing numbers of mechanisms. Given that the number of mechanisms could be reliably inferred from data for all variables, the meta-causal state of a graph could be identified as all unique combinations of mechanisms that are jointly active at a certain point in time. To recover the full set of meta-causal states, one needs to be able to simultaneously estimate the triple of active parents for every mechanism, the mechanisms parameterizations, and the resulting meta-state assignment of all data points for every meta-state of the system. All of the three components can vary between each meta-causal state (e.g. edges vanishing or possibly switching direction, thus altering the parent sets). Note that whenever any of the three components is known, the problem becomes rather trivial. However, without making any additional assumptions on the model or data and given the results on the already challenging task of identifying mechanisms in the bivariate case, the extension to a unsupervised full-fledged meta-causal state discovery is not obvious to us at the time of writing and we leave it to future work to come up with a feasible algorithm.
@@ 4.3 A Meta-State Analysis on Stress-Induced Fatigue
Figure 2: Mechanistic Decomposition as Meta-Causal States. (left) Effect of the stress level on itself (orange) plotted against the identity (blue). Once a threshold of $0.5$ is reached, the function switches behavior from self-suppressing self-reinforcing. (center) Contribution of the stress level mechanism for varying external stressors. Red arrows indicate a self-reinforcing effect, while green arrows indicate suppression. The gray area indicates the system without external stressors. Although all states are governed by the same set of structural equations, our meta-causal analysis identifies mechanistic differences and assigns different behavioral states to the different starting conditions. (right) The standard SCM gets decomposed into different meta-causal states. While the graph adjacency remains the same, the identification function detect the different modes of the causal mechanism. (Best viewed in color.)
@@
We use an identification function to distinguish between two different modes of operation of a causal mechanism. In particular, we are interested whether the inherent dynamics of the internal stress level are self-reinforcing or self-suppressing. For easier analysis of the system we decompose the dynamics of the internal stress variable into a ‘decayed stress’ $d$ and ‘resulting stress’ $s$ computation. The first term are the previous stress levels decayed over time with external factors added. The resulting stress is then the output of a Sigmoidal function (Fig. 2, left) that either reinforces or suppresses the value (Fig. 2,center). The structural equations are defined as follows and we assume that all values lie in the interval of $[0,1]$:
$f_{d}$ $:=0.95~{}\text{clip}_{[0,1]}(s^{\prime}+0.5\times\text{externalStress})$
$f_{s}$ $:=1.01(\frac{1}{1+\text{exp}({-15x+7.5})}-0.5)+0.5$
@@
While being rather simple in setup, this example serves as a good demonstration on how dynamical systems induce meta-causal states that exhibit qualitative different behavior although being governed by the same set of underlying structural equations.
Role of Latent Conditioning. A key takeaway in this example is that the current meta-causal state persists due to the current stress levels and system dynamics and not due to a chance in the structural equations. As a consequence, the system can not solely be expressed via some external conditioning factors. In contrast to other examples where system dynamics where purely due to the meta-causal state, here the variable values themselves play a vital role in the overall system dynamics. As a result, the stressed state of a person would persist even when the initiating external stressors disappear. Intervening on the meta-causal state of the system is now ill-defined, as both positive and negative reinforcing effects are governed by the same equation. Thus, creating a disparity between the intervened meta-causal state and the systems’ identified functional behavior.

### v2

@@ 4 Applications
In this section, we discuss several applications of the meta-causal formalism. First, we revisit the motivational example and consider how meta-causal models can be used to attribute responsibility. Next, we identify the presence of multiple mechanisms for the bivariate case from sets of unlabeled data. Finally, we analyze the emergence of meta-causal states from a dynamical system, highlighting that meta-causal states are more expressive than simple conditionings of the adjacency matrix. Code is made available at https://github.com/MoritzWillig/metaCausalModels.
@@ 4.1 Attributing Responsibility
Note that in this scenario, a classic counterfactual consideration, $A_{X}^{A_{\operatorname{\bm{\pi}}}=\text{standing still}}-A_{X}^{A_{\operatorname{\bm{\pi}}}=\text{following}}$, would also have inferred an effect of $A_{\operatorname{\bm{\pi}}}$ on $A_{X}$ from a purely value-based perspective. Attributing effects via observed changes in variable values is a valid approach, but it fails to explain preventative mechanisms. For example, consider the simple scenario where two locks prevent a door from opening. Classical counterfactual analysis would attribute zero effect to the opening of either lock. A meta-causal perspective would reveal that the opening of the first lock already changes the meta-causal state of the model by removing its causal mechanisms that condition the state of the door. In a sense, our meta-causal causal perspective is similar to that of the actual causality framework (Halpern, 2016; Chockler & Halpern, 2004). However, actual causality operates at the ‘actual’, i.e. value-based, level and does not take the mechanistic meta-causal view into perspective. Meta-causal models already allow us to reason about causal effects after opening the first lock due to a change in the causal graph, without us having to consider the effects of opening the second lock at some point to reach that conclusion.
@@ 4.2 Discovering Meta-Causal States in the Bivariate Case
Our goal in this experiment is to recover the number of meta-causal states $K\in[1..4]$ from data that exists between two variables $X,Y$ that are directly connected by a linear equation with added noise. We assume that each meta-causal state gives rise to a different linear equation $f_{k}:=\alpha_{k}X+\beta_{k}+\mathcal{N},k\in\mathbb{N}$, where $\alpha_{k},\beta_{k}$ are the slope and intercept of the respective mechanism and $\mathcal{N}$ is a zero-centered, symmetric, and quasiconvex noise distribution [^1] Implying unimodality and monotonic decreasing from zero allows us to distinguish the noise mean and intercept and to recover the parameters from a simple linear regression. Without loss of generality, we apply Laplacian noise, for which an L1-regression can estimate the true parameters of the linear equations (Hoyer et al., 2008). The causal direction of the mechanism is randomly chosen between different meta-causal states. The exact sampling parameters and plots of the resulting distributions are described in Sec. G (and plots of sample distributions are shown in Fig. 3 in the Appendix). In general, this scenario corresponds to the setting described above of inferring the values of a latent variable $Z$ (with $K=|Z|$) that indicate a particular meta-causal state of the system. Our goal is to recover the number of parameterizations of the causal mechanisms, and as a consequence to be able to reason about which points were generated by which mechanism.
@@
Approach. The problem we are trying to solve is twofold : first, we are initially unaware of the underlying meta-causal state $t$ that generated a particular data point $(x_{i},y_{i})$, which prevents us from estimating the parameterization $(\alpha_{k},\beta_{k})$ of the mechanism. Conversely, our lack of knowledge about the mechanism parameterizations $(\alpha_{k},\beta_{k})_{k\in[1..K]}$ prevents us from assigning class probabilities to the individual data points. Since neither the state assignment nor the mechanism parameterizations are initially known, we perform an Expectation-Maximization (EM; Dempster et al. (1977)) procedure to iteratively estimate and assign the observed data points to the discovered meta-causal state parameterizations. Due to the local convergence properties of the EM algorithm, we further embed it into a locally optimized local random sample consensus approach (LO-RANSAC; Fischler & Bolles (1981); Chum et al. (2003)). RANSAC approaches repeatedly sample initial parameter configurations to avoid local minima, and successively perform several steps of local optimization - here the EM algorithm - to regress the true parameters of the mechanism.
@@
Assuming for the moment that the correct number of mechanisms $k^{*}$ has been chosen, we assume that the EM algorithm is able to regress the parameters of the mechanisms, $\alpha_{k}^{*},\beta_{k}^{*}$, whenever there exists a pair of points for each of the mechanisms, where both points of the pair are samples generated by that particular mechanism. The chances of sampling such an initial configuration decrease rapidly with an increasing number of mechanisms (e.g. $0.036\%$ probability for $k=4$ and equal class probabilities). Furthermore, we assume that the sampling probabilities of the individual mechanisms in the data can deviate from the mean by up to a certain factor $d$. In our experiments, we consider setups with $d\in\{0.0,0.1,0.2\}$. Given the number of classes and the maximum sample deviation of the mechanisms, one can compute an upper bound on the number of resamples required to have a $95\%$ chance of drawing at least one valid initialization. The bound is maximized by assigning the first half of the classes the maximum deviation probability $P_{\text{k-max}}=(1+d)/k$ and the other half the minimum deviation probability $P_{\text{k-min}}=(1-d)/k$. We provide the formulas for the upper bound estimation and in Sec. E in the Appendix and provide the calculated required resample counts in Table 2. In the worst case, for a scenario with $k=4$ mechanisms and $d=0.2$ maximum class probability deviation, nearly $14,900$ restarts of the EM algorithm are required, drastically increasing the potential runtime.
@@
Analysis Type 1 Mechanism 2 Mech. 3 Mech. 4 Mech.
Max. Class Theoretical 1 23 363 8,179
Deviation = 0.0 Empirical 2 8 24 173
Max. Class Theoretical 1 26 429 10,659
Deviation = 0.1 Empirical 2 8 25 177
Max. Class Theoretical 1 30 526 14,859
Deviation = 0.2 Empirical 2 8 26 177
Table 2: Estimated Number of Required Resamples for Obtaining a 95% Convergence Rate with the LO-RANSAC Algorithm per Number of Mechanisms and Maximum Class Deviation. The empirical observed convergence rates of the EM algorithm drastically reduce the theoretical derived bound of required samples. This reveals that convergence assumptions where chosen to be quite conservative, and attests a good fit of the EM algorithm for regressing mechanism parameters.
@@
The EM algorithms return the estimated parameters $\alpha^{\prime}_{k},\beta^{\prime}_{k}$, and the mean standard deviation $b^{\prime}_{k}$ of the Laplacian [^2] Since mechanisms can go in both directions, $X\rightarrow Y$ and $Y\rightarrow X$ (cyclic relations are not considered), we repeat the regression for both directions and use an Anderson-Darling test (Anderson & Darling, 1952) on the residual to test which of the distributions more closely resembles a Laplacian distribution at each step.. This allows one to compute the class densities $k(x_{i},y_{i};\alpha^{\prime}_{k},\beta^{\prime}_{k},b^{\prime}_{k})$ for a pair of values $(x_{i},y_{i})$. Since the assignment of mechanisms to a data point may be ambiguous due to the overlap in the estimated PDFs, we normalize the density values of all mechanisms per data point and consider only those points for which the probability of the dominant mechanism is $0.4$ higher than the second class: $(\mathbf{x}^{j},\mathbf{y}^{j}):=((x_{i},y_{i})|\#1=j;f^{\#2}_{i}<P^{\#1}_{i}\times(1-0.4))$, where #$\#n$ indicates the class with the n-th highest density value. Finally, the residuals $\mathbf{y}^{j}-f^{\prime j}_{k}(\mathbf{x}^{j},\mathbf{y}^{j};\alpha^{\prime}_{k},\beta^{\prime}_{k},b^{\prime}_{k})$ are computed for all data points $(\mathbf{x}^{j},\mathbf{x}^{j})$ assigned to a particular predicted mechanism $f^{\prime j}_{k}$. We choose the parameterization that best fits the data for each $k\in[1..4]$ and, make use of the Anderson-Darling test (Anderson & Darling, 1952) to test the empirical distribution function of the residuals against the Laplacian distribution with an $\alpha=0.95$ using significance values estimated by Chen (2002). If all residual distributions of all mechanisms for a given $k^{\prime}$ pass the Anderson-Darling test, we choose that number as our predicted number of mechanisms. If the algorithm finds that none of the $k\in[1..4]$ setups pass the test, we refrain from making a decision. We provide the pseudo code for our method in Algorithm. 1 in the Appendix.
@@
Evaluation and Results. We evaluate our approach over all $k\in[1..4]$ by generating 100 different datasets for every particular number of mechanisms. For every dataset we sample 500 data points from each mechanism $(x^{k}_{i},y^{k}_{i})=f_{k}(\alpha_{k}x^{k}_{i}+\beta_{k}+l_{i})$ where $l_{i}\sim\mathcal{L}(0,b_{k})$, using the same sampling method as before (c.f. Appendix G). Finally, the algorithm recovers the number of mechanisms.
@@
We compare theoretically computed and empirical in convergence results in Table 2. In practice, we observe that the convergence of the EM algorithm is more favorable than estimated, reducing from $23$ to $8$ required examples for the simple case of $k=2,d=0.0$, and requiring up to 83-times fewer samples for the most challenging setup of $k=4,d=0.2$, reducing from a theoretical of 14,859 to an empirical estimate of 177 samples. The actual convergence probabilities and the formula for deriving to sample counts are given in Table 3 and Sec. F of the Appendix.
@@
Extension to Meta-Causal State Discovery on Graphs. Our results indicate that identifying meta-causal mechanisms even in the bivariate case comes with an increasing number of uncertainty when it comes to increasing numbers of mechanisms. Given that the number of mechanisms could be reliably inferred from data for all variables, the meta-causal states could be identified as all unique combinations of mechanisms that are jointly active at a certain point in time. To recover the full set of meta-causal states, one needs to be able to simultaneously estimate the triple of active parents for every mechanism, the mechanisms parameterizations, and the resulting meta-state assignment of all data points for every meta-state of the system. All of the three components can vary between each meta-causal state (e.g. edges vanishing or possibly switching direction, thus altering the parent sets). Note that whenever any of the three components is known, the problem becomes rather trivial. However, without making any additional assumptions on the model or data and given the results on the already challenging task of identifying mechanisms in the bivariate case, the extension to a unsupervised full-fledged meta-causal state discovery is not obvious to us at the time of writing and we leave it to future work to come up with a feasible algorithm.
@@ 4.3 A Meta-State Analysis on Stress-Induced Fatigue
Figure 2: Mechanistic Decomposition as Meta-Causal States. (left) The effect of the stress level on itself (orange) plotted against the identity (blue; corresponding to a non-self-reinforcing effect). Once a certain threshold is reached, the function switches its behavior from self-suppressing to a self-reinforcing effect. (center) Contribution of the stress level mechanism for varying external stressors. Red arrows indicate a self-reinforcing effect, while green arrows indicate a suppressive effect. The gray area highlights the system configuration without external stressors. Although all states are governed by the same structural equations, our meta-causal analysis identifies the mechanistic difference and decomposes the corresponding initial conditions into different meta-causal states. (right) The standard SCM gets decomposed into different meta-causal states. While the graph adjacency remains the same, the different starting conditions identify different behavioral types of the causal mechanism. (Best viewed in color.)
@@
While being rather simple in setup, the example serves as a good demonstration on how dynamical systems induce meta-causal states that exhibit qualitative different behavior, while employing the same set of underlying structural equations. As such, the inherent behavior of the system is not only due to a conditioning external factor. We use an identification function to distinguish between two different modes of operation of a causal mechanism. In particular, we are interested whether the inherent dynamics of the internal stress level are self-reinforcing or self-suppressing. For easier analysis of the system we decompose the dynamics of the internal stress variable into a ‘decayed stress’ $d$ and ‘resulting stress’ $s$ computation. The first term are the previous stress levels decayed over time with external factors added. The resulting stress is then the output of a Sigmoidal function (Fig. 2, left) that either reinforces or suppresses the value (Fig. 2,center). The structural equations are defined as follows and we assume that all values lie in the interval of $[0,1]$:
[…]
@@
Role of Latent Conditioning. A key takeaway of this example is that the current meta-state persists due to the inherent stress level and dynamics of the system. In contrast to other examples where system dynamics where purely due to the meta-causal state, variable values play a role in the overall system dynamics. As a result, the stressed state of a person would persist even when the initiating external stressors disappear. Intervening on the meta-causal state of the system is now ill-defined, as both positive and negative reinforcing effects are governed by the same equation. Thus, creating a disparity between the intervened meta-causal state and the systems’ identified functional behavior.